# Efficient Unsupervised Domain Adaptation Regression for Spatial-Temporal Air Quality Sensor Fusion

### v1

@@ Abstract
The deployment of affordable Internet of Things (IoT) sensors for air pollution monitoring has increased in recent years due to their scalability and cost-effectiveness. However, accurately calibrating these sensors in uncontrolled environments remains a significant challenge. While expensive reference sensors can provide accurate ground truth data, they are often deployed on a limited scale due to high costs, leading to a scarcity of labeled data. In diverse urban environments, data distributions constantly shift due to varying factors such as traffic patterns, industrial activities, and weather conditions, which impact sensor readings. Consequently, traditional machine learning models – despite their increasing deployment for environmental sensor calibration– often struggle to provide reliable pollutant measurements across different locations due to domain shifts. To address these challenges, we propose a novel unsupervised domain adaptation (UDA) method specifically tailored for regression tasks on graph-structured data. Our approach leverages Graph Neural Networks (GNNs) to model the relationships between sensors. To effectively capture critical spatial-temporal interactions, we incorporate spatial-temporal graph neural networks (STGNNs), which extend GNNs by incorporating temporal dynamics. To handle the resulting larger embeddings, we propose a domain adaptation method using a closed-form solution inspired by the Tikhonov-regularized least-squares problem. This method leverages Cholesky decomposition and power iteration to align the subspaces between source and target domains. By aligning these subspaces, our approach allows low-cost IoT sensors to learn calibration parameters from expensive reference sensors. This facilitates reliable pollutant measurements in new locations without the need for additional costly equipment. Experimental evaluations on real-world datasets show that our method achieves competitive performance while significantly reducing training time, making it practical for large-scale air pollution monitoring in urban environments.
keywords:
@@
Unsupervised domain adaptation , Graph neural networks , Multisensor fusion , Internet of Things , Air quality , Smart cities
journal: ArXiv
[…]
[inst1]organization=Intelligent Maintenance and Operations Systems (IMOS) Lab, EPFL, Switzerland
1 Introduction
The increase in air pollution, driven by growing urban population densities, escalating vehicular traffic, and expanding industrial emissions, poses a significant threat to both public and environmental health. Various pollutants such as carbon monoxide (CO), nitrogen dioxide (NO2), and tropospheric ozone (O3) are particularly hazardous [1]. Therefore, it is essential to establish reliable air quality monitoring systems that focus on different contaminants to track environmental progress, provide warnings to inhabitants, and support informed decision-making for public health and safety. Such monitoring contributes to the overall objectives of developing environmentally friendly and livable urban environments, characteristics of smart and sustainable cities [2]. These systems are also essential for safeguarding public health and contributing to the effective management of pollutants.
The advent of the Internet of Things (IoT) has played a crucial role in advancing smart and sustainable urban environments [3]. IoT-based technologies have made it possible to collect real-time data from various sensors installed and distributed across different locations [4]. However, despite these advancements, concerns persist regarding the accuracy of low-cost sensors integrated into IoT-based solutions [5, 6]. Manufacturers typically do not calibrate environmental sensors, and if they do, the calibration is not performed under the actual environmental conditions where the sensors will operate. This can negatively impact the performance of the sensor network [7]. To address this challenge, many air quality monitoring platforms incorporate inexpensive temperature and humidity sensors alongside gas sensors to enable reliable calibration [8]. Nevertheless, implementing low-cost sensor networks presents specific challenges such as sensor drift, interference from environmental factors, and discrepancies in sensor calibration, which can lead to inaccurate data readings. These sensor networks often struggle with limited sensitivity, making it difficult to measure specific contaminants precisely. Moreover, sensor network maintenance and data validation can be complex, as individual sensor reliability may vary, affecting overall data quality and consistency [9]. These challenges can significantly hinder the utilization of such sensor networks in regulatory pollutant measurements, necessitating the use of expensive and high-precision measuring devices as references for calibration. One potential strategy to enhance the reliability of low-cost sensor networks involves incorporating hardware redundancy across the sensor network. Additionally, employing multisensor fusion methods could enable the utilization of all low-cost sensors to achieve highly reliable measurements for various pollutants [10].
@@ 1 Introduction
Traditional machine learning algorithms such as support vector regression (SVR), random forest (RF), and k-nearest neighbors (KNN) have been shown to be feasible options for multisensor fusion and calibration in IoT-based air pollution monitoring platforms [7]. Despite their fairly good performance, these traditional machine learning models still face challenges in achieving reliable estimation of air pollutants, limiting their deployment reliability. One limitation of these models is their dependency on extensive feature engineering before data input. For instance, air quality sensor networks often experience multicollinearity, where multiple sensors provide highly correlated information. This redundancy not only diminishes model performance but also leads to inefficient training processes. Effective feature engineering in such contexts requires substantial domain knowledge and can be highly specialized, presenting challenges in developing reliable machine learning models. Moreover, these models typically fail to account for the spatial and temporal interactions among sensor measurements that are particularly relevant in air quality sensor networks. Although temporal pattern-based denoising can improve performance as a post-processing step following fusion and calibration, as indicated in [11], there remains a critical need for a model that can jointly capture both spatial and temporal features in a comprehensive, end-to-end approach.
Spatial-temporal graph neural networks (STGNNs) have emerged as powerful tools in various applications due to their ability to capture complex spatial-temporal relationships within data [12]. This versatility has been demonstrated in different fields such as traffic data analysis [13, 14, 15], industrial IoT [16, 17], and evolving social networks [18]. Leveraging their capability to capture spatial-temporal interactions among sensor measurements, a recent study applied STGNNs for multisensor fusion and calibration in IoT air pollution monitoring platforms [19]. This study showed that this architecture can achieve promising results compared to other data-driven methods that did not leverage the graph structure of air quality sensor networks.
@@
STGNNs typically perform well when the testing data distribution aligns with the training data distribution. However, in real-world scenarios, especially in diverse urban environments, data distributions are prone to change over time. This shift can result in performance degradation. This occurs because most current data-driven models assume that the statistical properties of data remain the same [20]. In the context of air quality monitoring, geographical variations due to factors such as weather conditions, pollution sources, and urban planning can cause significant differences in data distributions. Consequently, models trained in one location may not perform when applied in another. A straightforward but expensive solution would be to collect new labeled data at each specific measurement location and train a separate model for each deployment site. This method requires high-precision instruments at each site for accurate calibration, which can be prohibitively costly. Therefore, an alternative approach is needed that uses available labeled data from a specific location and adapts the model to different locations without relying on high-precision sensors at each site.
@@
To address this challenge, unsupervised domain adaptation (UDA) shows promise in transferring knowledge from a labeled source domain to a related but unlabeled target domain [21]. UDA would allows for the utilization of only a single high-precision sensor in the source domain, eliminating the need for labeled data in the target domain. As a result, the need to deploy and manage multiple high-precision sensors across different environments is eliminated, reducing overall costs.
@@
Recently, novel UDA methods have emerged with a focus on regression tasks. Notably, Representation Subspace Distance for Domain Adaptation Regression (RSD) [29] aims to align orthogonal bases while preserving the feature scale. The RSD-based approach has shown improved results, highlighting the need for UDA methods specifically designed for regression tasks.
Alternatively, the DARE-GRAM [30] method approaches UDA for regression from a regressor’s perspective. Motivated by the closed-form least square solution, this method proposes aligning the pseudo-inverse of the Gram matrix instead of directly aligning the features. DARE-GRAM, which focuses on scale and angle alignment based on the Gram matrix, results in better-calibrated features for the regressor and exhibits less sensitivity to batch size compared to RSD.
Although these two subspace-based approaches [29, 30] provide state-of-the-art results for UDA in regression tasks, each training iteration requires computationally intensive singular value decomposition (SVD), which can be computationally too expensive for large-scale sensor networks.
@@
Building on this foundation, the primary goal of this paper is to develop a robust methodology for estimating pollutant levels using low-cost sensors in various urban environments. Our approach leverages the proposed TikUDA method to bridge the gap between two domains: training on labeled data from low-cost sensors calibrated with a high-precision sensor in one location (the source domain) and transferring that knowledge to a different location (target domain) using unlabeled data from low-cost sensors.
To capture both spatial and temporal dependencies in sensor measurements, we propose using spatial-temporal graph neural networks (STGNNs) within IoT air quality sensor networks. The model first encodes input data into a hidden space with a linear encoder. Temporal dependencies are then extracted using a Gated Recurrent Unit (GRU), followed by spatial pattern extraction using a Graph Attention Network (GAT). To ensure the model generalizes well to the target location, we apply our TikUDA method to align the feature representations from both the source and target domains.
Our experimental results, based on real-world datasets, demonstrate the effectiveness of this approach in estimating air pollutants and improving generalization to new target locations using low-cost sensors. This significantly reduces the dependency on numerous high-precision reference sensors, thus lowering both cost and system complexity. To the best of our knowledge, this study is the first to apply UDA in the context of multisensor fusion and calibration for IoT air quality sensor networks, providing a scalable and efficient solution for improving air quality monitoring in smart and sustainable cities.
@@
    We evaluate the proposed method on real-world IoT air quality sensor network datasets, showcasing its practical applicability and superior performance in real-world applications.
@@
The remainder of the paper is organized as follows: Section 2 provides an in-depth introduction to the relevant background in multisensor fusion and unsupervised domain adaptation for regression tasks. In Section 3, we present the proposed spatial-temporal feature extractor along with our proposed TikUDA method for unsupervised domain adaptation. Sections 4 and 5 detail the experimental setup and results, illustrating the efficacy of our approach compared to various baseline and state-of-the-art methods. Finally, Section 6 summarizes the findings and suggests possible directions for future research.
2 Problem Definition
2.1 Multisensor Fusion and Calibration
Multisensor data fusion and calibration involve the integration of data from multiple sensors, denoted as $\{s_{1},...,s_{n}\}$, within an IoT air quality sensor network. These sensors measure various environmental parameters such as pollutants, humidity, and temperature. The goal is to produce highly reliable estimates of pollutant levels, including ozone and nitrogen dioxide values, by aggregating data from a set of sensors that include not only unreliable gas measurement sensors but also measure temperature and humidity as illustrated in Figure 1.
Figure 1: Diagram for Multisensor Fusion in IoT Air Quality Monitoring.
Given the spatial-temporal characteristics of urban data, several studies have integrated GNNs with different learning strategies to capture dynamics in both space and time [31]. Building on this research, we propose using a Spatial-Temporal Graph Neural Network (STGNN) model for IoT air quality sensor prediction. STGNNs are designed to capture both spatial and temporal relationships within data by leveraging the spatial interactions among sensors and the temporal patterns in their measurements.
However, relying solely on direct predictions from low-cost IoT air quality sensors is insufficient for training an accurate multisensor fusion model in new locations. In this paper, we propose leveraging labeled data from low-cost IoT air quality sensor networks deployed in a specific urban area. These data are strategically collected from sites near high-precision sensors used by governmental agencies, serving as ground truth labels in the source domain. Additionally, we use unlabeled data obtained from low-cost IoT air quality sensors in new locations (target domain), which are sufficiently distant from the source location to experience a domain shift. To counteract the performance degradation caused by this shift and to deploy the model in the new location without high-precision sensors, we employ unsupervised domain adaptation (UDA). This approach mitigates the impact of domain shift and ensures accurate predictions of pollutant levels. We assume that the IoT air quality sensor networks in both locations maintain an identical number of sensors.
2.2 Unsupervised Domain Adaptation
@@
[…]
where $\mathbf{y}^{tgt}$ denotes the ground truth label for the target input feature $\mathbf{x}^{tgt}$ and is not available during training.
@@
3 Methodology
In this work, we propose a Tikhonov-regularized least squares-based unsupervised domain adaptation (TikUDA) method. This method improves over previous methods, as detailed in Section 3.1. Furthermore, we introduce a Spatial-Temporal Graph Neural Network (STGNN) designed to effectively capture both spatial and temporal interactions in multisensor fusion tasks. This model leverages the spatial correlations between different sensor nodes and temporal dependencies within their measurements, providing a promising framework for integrating and analyzing information from multiple sensors over time, as detailed in Section 3.2. In Section 3.3, we provide a comprehensive overview and summary of the overall framework.
3.1 TikUDA
To alleviate the ill-conditioned problem of the Gram matrix, we propose using the closed-form solution of Ridge regression – a specific case of Tikhonov-regularized least-squares – instead of using the closed-form solution of OLS. Specifically, we apply the Tikhonov matrix $\mathbf{Z}^{\mathsf{T}}\mathbf{Z}+\alpha\mathbf{I}$, which ensures the Gram matrix is fully ranked by adding the identity matrix. This addition guarantees the smallest eigenvalue is never zero but it depends on $\alpha$. This eliminates the need to compute the SVD to threshold degenerate spaces (small eigenvalues) and allows for direct inverse matrix decomposition. Following DARE-GRAM methodology [30], we propose aligning the subspaces of the source and target domains using the angle between the columns of the inverse Tikhonov matrix, which are ordered based on the importance of their eigenvalue. To ensure consistent scaling between the source and target feature representations, we align their eigenvalues, addressing the scale sensitivity in regression models as highlighted in [29].

### v2

# Efficient Unsupervised Domain Adaptation Regression for Spatial-Temporal Sensor Fusion

@@ Abstract
The growing deployment of low-cost, distributed sensor networks in environmental and biomedical domains has enabled continuous, large-scale health monitoring. However, these systems often face challenges related to degraded data quality caused by sensor drift, noise, and insufficient calibration – factors that limit their reliability in real-world applications. Traditional machine learning methods for sensor fusion and calibration rely on extensive feature engineering and struggle to capture spatial-temporal dependencies or adapt to distribution shifts across varying deployment conditions. To address these challenges, we propose a novel unsupervised domain adaptation (UDA) method tailored for regression tasks. Our proposed method integrates effectively with Spatial-Temporal Graph Neural Networks and leverages the alignment of perturbed inverse Gram matrices between source and target domains, drawing inspiration from Tikhonov regularization. This approach enables scalable and efficient domain adaptation without requiring labeled data in the target domain. We validate our novel method on real-world datasets from two distinct applications: air quality monitoring and EEG signal reconstruction. Our method achieves state-of-the-art performance which paves the way for more robust and transferable sensor fusion models in both environmental and physiological contexts. Our code is available at https://github.com/EPFL-IMOS/TikUDA.
Index Terms:
@@
Unsupervised domain adaptation, Graph neural networks, Multisensor fusion, Internet of Things, Air quality, Healthcare
I Introduction
The growing deployment of low-cost, distributed sensor networks has fundamentally reshaped how real-time data is collected and utilized across a wide range of applications – from urban air quality monitoring to physiological sensing. In environmental contexts, sensor arrays are deployed to measure key pollutants such as carbon monoxide (CO), nitrogen dioxide (NO2), and tropospheric ozone (O3), to enabling data-informed public health interventions, exposure alerts, and evidence-based urban planning [1, 2]. In biomedical domains, wearable sensors like electroencephalography (EEG) continuously record physiological signals to support clinical diagnostics, brain–computer interfaces, and patient monitoring. Together, these systems underpin efforts to protect both public health and the environment by enabling timely, data-driven decision-making across very different but equally critical domains.
The adoption of Internet of Things (IoT) architectures has enabled the large-scale deployment of low-cost sensor networks for real-time monitoring in both environmental and biomedical domains [3, 4]. However, these low-cost sensing platforms suffer from several interrelated accuracy issues. In air quality networks, gas sensors are rarely calibrated under true field conditions, resulting in biases and drift as ambient temperature, humidity, and pollutant concentrations fluctuate [5, 6, 7]. In physiological monitoring, wearable electrodes and inexpensive biosensors exhibit signal degradation due to electrode displacement, motion artifacts, and pronounced inter-subject variability, which together induce degradation of the recorded data. To mitigate these effects, environmental systems often incorporate auxiliary temperature and humidity sensors, while biomedical setups may co-register signals with higher-precision reference instruments or additional modalities [8]. Although these measures can improve raw measurement fidelity, they introduce significant hardware complexity, increase maintenance demands, and do not fully eliminate cross-sensitivity or calibration inconsistencies. Consequently, both regulatory applications (e.g., pollutant compliance monitoring) and clinical use cases (e.g., patient diagnostics) remain constrained by limited sensitivity and variable data quality [9]. Two principal strategies have emerged to address these limitations: hardware redundancy—deploying overlapping low-cost sensors to average out individual device errors—and advanced multisensor fusion algorithms that jointly process multiple data streams to produce robust, high-precision estimates without exclusive reliance on expensive reference equipment [10].
@@ I Introduction
Traditional machine learning algorithms such as support vector regression (SVR), random forest (RF), and k-nearest neighbors (KNN) have been shown to be feasible options for multisensor fusion and calibration in IoT-based monitoring platforms [7]. Despite their fairly good performance, these traditional machine learning models still face challenges in achieving reliable estimation of air pollutants, limiting their deployment reliability. One limitation of these models is their dependency on extensive feature engineering before data input. For instance, sensor networks often experience multicollinearity, where multiple sensors provide highly correlated information. This redundancy not only diminishes model performance but also leads to inefficient training processes. Effective feature engineering in such cases requires substantial domain knowledge and can be highly specialized, presenting challenges in developing reliable machine learning models. Moreover, these models typically fail to account for the spatial and temporal interactions among sensor measurements that are particularly relevant in environmental or biomedical sensor networks. Although temporal pattern-based denoising can improve performance as a post-processing step following fusion and calibration, as indicated in [11], there remains a critical need for a model that can jointly capture both spatial and temporal features in a comprehensive, end-to-end approach.
@@
Spatial-temporal graph neural networks (STGNNs) have emerged as powerful tools in various applications due to their ability to capture complex spatial-temporal relationships within data [12]. This versatility has been demonstrated in different fields such as traffic data analysis [13, 14, 15], industrial IoT [16, 17], and evolving social networks [18]. Leveraging their capability to capture spatial-temporal interactions among sensor measurements, a recent study applied STGNNs for multisensor fusion and calibration in IoT air pollution monitoring platforms [19]. This study showed that this architecture can achieve promising results compared to other data-driven methods that did not leverage the graph structure of air quality sensor networks. Similarly, STGNNs have been shown to effectively model the spatial relationships among EEG channels and the temporal dynamics of neural activity, leading to improved classification performance in EEG‐based tasks, as highlighted in a recent survey of STGNN applications for EEG signal processing [20]. STGNNs typically perform well when the testing data distribution aligns with the training data distribution. However, in real-world scenarios, especially in diverse urban environments and different subjects, data distributions are prone to change over time. This shift can result in performance degradation. This occurs because most current data-driven models assume that the statistical properties of data remain the same [21]. In air quality monitoring platforms, geographical variations due to factors such as weather conditions, pollution sources, and urban planning can cause significant differences in data distributions. Similarly, in EEG applications, distribution shifts arise from inter‐subject variability, differences in electrode placement, individual neurophysiology, and hardware configurations, leading to degraded model performance when transferring across users or recording setups. Consequently, models trained in one location or subject may not perform when applied in another. A straightforward but expensive solution would be to collect new labeled data at each specific subject or measurement location and train a separate model for each setting. However, this approach depends on deploying high-precision reference equipment such as regulatory-grade gas analyzers for each air quality station or high density EEG systems for every individual to achieve accurate calibration, which is prohibitively expensive. Consequently, there is a critical need for methods that can leverage labeled data from a source domain and adapt models to new target domains whether across different geographic locations or across subjects without requiring costly reference instruments at each deployment.
@@
To address this challenge, unsupervised domain adaptation (UDA) shows promise in transferring knowledge from a labeled source domain to a related but unlabeled target domain [22]. UDA would allow for the utilization of only a single high-precision sensor in the source domain, eliminating the need for labeled data in the target domain. As a result, the need to deploy and manage multiple high-precision sensors across different environments or subjects is eliminated, reducing overall costs.
@@
Recently, novel UDA methods have emerged with a focus on regression tasks. Notably, Representation Subspace Distance for Domain Adaptation Regression (RSD) [30] aims to align orthogonal bases while preserving the feature scale. The RSD-based approach has shown improved results, highlighting the need for UDA methods specifically designed for regression tasks. Alternatively, the DARE-GRAM [31] method approaches UDA for regression from a regressor’s perspective. Motivated by the closed-form least square solution, this method proposes aligning the pseudo-inverse of the Gram matrix instead of directly aligning the features. DARE-GRAM, which focuses on scale and angle alignment based on the Gram matrix, results in better-calibrated features for the regressor and exhibits less sensitivity to batch size compared to RSD. Although these two subspace-based approaches [30, 31] provide state-of-the-art results for UDA in regression tasks, each training iteration requires computationally intensive singular value decomposition (SVD), which can be computationally too expensive for large-scale sensor networks.
@@
Building on this foundation, the primary goal of this paper is to develop a robust methodology for multisensor fusion and regression across both environmental and physiological domains. Our approach leverages the proposed TikUDA method to bridge the gap between two domains: training on labeled data from low-cost sensors calibrated with a high-precision sensor in one location or subject (the source domain) and transferring that knowledge to a different location or subject (target domain) using unlabeled data from low-cost sensors.
To capture both spatial and temporal dependencies in sensor measurements, we propose using spatial-temporal graph neural networks (STGNNs) within sensor networks. The model first encodes input data into a hidden space with a linear encoder. Temporal dependencies are then extracted using a Gated Recurrent Unit (GRU), followed by spatial pattern extraction using a Graph Attention Network (GAT). To ensure the model generalizes well to the target location, we apply our TikUDA method to align the feature representations from both the source and target domains.
Using real-world datasets, we evaluate our framework on two different tasks. First, in air pollutant estimation, our method accurately predicts concentrations of pollutant at unseen locations using only low-cost sensor data, thereby reducing reliance on high-precision reference instruments and lowering both deployment cost and system complexity. Second, in EEG signal processing, the approach successfully reconstructs a missing channel from neighboring electrodes, enhancing signal fidelity without additional hardware.
@@
    We evaluate the proposed method on two real-world sensor networks, demonstrating its practical applicability and superior performance in real-world applications.
@@
The remainder of the paper is organized as follows: Section II provides an in-depth introduction to the relevant background in multisensor fusion and unsupervised domain adaptation for regression tasks. In Section III, we present the proposed spatial-temporal feature extractor along with our proposed TikUDA method for unsupervised domain adaptation. Sections IV and V detail the experimental setup and results, illustrating the efficacy of our approach compared to various baseline and state-of-the-art methods. Finally, Section VI summarizes the findings and suggests possible directions for future research.
II Problem Definition
II-A Multisensor Fusion and Calibration
Multisensor data fusion and calibration involve the integration of data from multiple sensors, denoted as $\{s_{1},...,s_{n}\}$. In air quality monitoring, these sensors record pollutants, temperature, and humidity; in EEG applications, electrodes capture neural signals across multiple scalp locations. The aim is to produce reliable estimates of target variables by aggregating multiple sensors, as illustrated in Figure 1.
(a) Multisensor fusion in IoT air quality monitoring to predict accurate quantities of various pollutants
(b) Sensor fusion for estimating missing EEG channel (Channel 2).
Figure 1: Overview of sensor fusion tasks across different applications. (a) Air quality (b) EEG signal processing
Given the spatial‐temporal nature of the sensor networks, recent work has applied STGNNs to model sensor interactions and temporal evolution [32]. STGNNs leverage the graph structure—connecting co‐located gas sensors or neighboring EEG electrodes—to jointly learn spatial dependencies and temporal patterns, outperforming methods that treat these aspects separately.
However, relying solely on direct predictions from low-cost IoT air-quality sensors or on simple interpolation of EEG signals is inadequate for accurate multisensor fusion when models are deployed in new environments or on new subjects. In our approach, we leverage labeled source-domain data from low-cost sensor networks co-located with high-precision reference monitors for air-quality monitoring and from complete EEG recordings collected in an initial subject cohort, while treating as the target domain the unlabeled low-cost sensor measurements obtained in different geographic locations and the incomplete EEG recordings with a missing channel in new participants. Although both source and target domains maintain the same number of sensors or electrodes, distributional shifts arise from local environmental factors or inter-subject variability. To address this, we proposed new domain adaptation methods to align spatial–temporal feature representations across domains. This framework enables accurate estimation of pollutant concentrations and robust reconstruction of missing EEG channels without requiring high-precision instruments or full electrode setups at every deployment.
II-B Unsupervised Domain Adaptation
@@
$\underset{\mathbf{\theta}}{\arg\min}\,\mathbb{E}\{\|f_{\mathbf{\theta}}(\mathbf{x}^{tgt})-y^{tgt}\|_{2}^{2}\},$ (1)
where ${y}^{tgt}$ denotes the ground truth label for the target input feature $\mathbf{x}^{tgt}$ and is not available during training.
@@
III Methodology
In this work, we propose a Tikhonov-regularized least squares-based unsupervised domain adaptation (TikUDA) method. This method improves over previous methods, as detailed in Section III-A. Furthermore, we introduce an STGNN designed to effectively capture both spatial and temporal interactions in multisensor fusion tasks. This model leverages the spatial correlations between different sensor nodes and temporal dependencies within their measurements, providing a promising framework for integrating and analyzing information from multiple sensors over time, as detailed in Section III-B. In Section III-C, we provide a comprehensive overview and summary of the overall framework.
III-A TikUDA
To alleviate the ill-conditioned problem of the Gram matrix, we propose using the closed-form solution of Ridge regression – a specific case of Tikhonov-regularized least-squares – instead of using the closed-form solution of OLS. Specifically, we apply the Tikhonov matrix $\mathbf{Z}^{\mathsf{T}}\mathbf{Z}+\alpha\mathbf{I}$, which ensures the matrix is fully ranked by adding the identity matrix. This addition guarantees the smallest eigenvalue is never zero but it depends on $\alpha$. This eliminates the need to compute the SVD to threshold degenerate spaces (small eigenvalues) and allows for direct inverse matrix decomposition. Following DARE-GRAM methodology [31], we propose aligning the subspaces of the source and target domains using the angle between the columns of the inverse Tikhonov matrix, which are ordered based on the importance of their eigenvalue. To ensure consistent scaling between the source and target feature representations, we align their eigenvalues, addressing the scale sensitivity in regression models as highlighted in [30]. Figure 2 illustrates the main motivation behind TikUDA. It presents two distributions representing features from the source and target domains. These feature representations differ in both scale and orientation, highlighting the challenges of aligning domains with different properties. The primary goal of this work is to align these distributions by adjusting both the scale and the angle of the subspaces they span. The alignment of Gram matrices provides a principled approach to mitigating domain shift by leveraging stable relational structures in representation space. The Gram matrix, capturing pairwise inner products, implicitly encodes the covariance-like structure of the data. While raw features are sensitive to domain-specific perturbations (e.g., sensor noise or background variations), these high-level relational geometries, including inter-vector distances and angles, tend to remain more invariant across domains. Mid-layer Gram matrices in deep networks further exhibit beneficial low-rank structure, concentrating meaningful variance into a few dominant directions while naturally filtering out noise and domain artifacts [33]. Aligning Gram matrices thus performs implicit manifold alignment, matching the intrinsic geometry of source and target feature spaces rather than exact feature values. This alignment enables more effective knowledge transfer between domains, which is crucial for improving performance in domain adaptation tasks.
Figure 2: Illustration of subspace-based domain adaptation. The source and target feature distributions differ in both scale and angle. The goal is to align the subspaces.